# Real-Time Per-Garment Virtual Try-On with Temporal Consistency for Loose-Fitting Garments

### v1

@@ Abstract.
Per-garment virtual try-on methods collect garment-specific datasets and train networks tailored to each garment to achieve superior results. However, these approaches often struggle with loose-fitting garments due to two key limitations: (1) They rely on human body semantic maps to align garments with the body, but these maps become unreliable when body contours are obscured by loose-fitting garments, resulting in degraded outcomes; (2) They train garment synthesis networks on a per-frame basis without utilizing temporal information, leading to noticeable jittering artifacts. To address these challenges, we propose a two-stage approach for robust semantic map estimation. First, we extract a garment-invariant representation from the raw input image. This representation is then passed through an auxiliary network to estimate the semantic map. This enhances the robustness of semantic map estimation under loose-fitting garments during garment-specific dataset generation. Furthermore, we introduce a recurrent garment synthesis framework that incorporates temporal dependencies to improve frame-to-frame coherence while maintaining real-time performance. We conducted qualitative and quantitative evaluations to demonstrate that our method outperforms existing approaches in both image quality and temporal coherence. Ablation studies further validate the effectiveness of the garment-invariant representation and the recurrent synthesis framework.
Virtual Try-On, Image Synthesis, Human-in-the-loop Machine Learning
copyright: noneconference: ; ; ccs: Computing methodologies Image processingccs: Human-centered computing Mixed / augmented realityccs: Computing methodologies Image-based rendering
@@
Figure 1. (a) Our virtual try-on system features an efficient inference pipeline that enables real-time performance. (b) Our method produces highly realistic $360^{\circ}$ virtual try-on results for loose-fitting garments with temporal consistency.
@@ 1. Introduction
Virtual try-on technology enables customers to try on garments without physical access, has recently garnered significant attention from researchers in today’s e-commerce landscape. Existing image-based virtual try-on methods (Jetchev and Bergmann, 2017; Dong et al., 2019; Choi et al., 2021; Lee et al., 2022; Fang et al., 2024; Karras et al., 2024) require only a 2D in-shop image of the target garment to generate virtual try-on results, making them highly accessible. However, these approaches are predominantly tailored for retailers aiming to produce promotional images featuring fashion models, rather than for end-users seeking personalized virtual try-on experiences. This focus is clearly reflected in the biased datasets used for training, which are heavily skewed toward tall and slim fashion models. As a result, these methods fail to generalize well to the diverse body types of average consumers.
Unlike general image-based methods, per-garment virtual try-on methods (Chong et al., 2021; Wu et al., 2024, 2025) focus on customer-oriented virtual try-on by collecting garment-specific datasets and training garment-specific neural networks. Early per-garment methods (Chong et al., 2021; Wu et al., 2024) require an expensive robotic mannequin for dataset collection, making their results non-reproducible. (Wu et al., 2025) lower the barrier to dataset collection significantly by using real human bodies for dataset collection. This approach relies on DensePose (Güler et al., 2018) for estimating human body semantic maps, which are crucial for aligning the human body and the synthesized garment in 2D image space. However, it struggles with loose-fitting garments, as DensePose results degrade significantly for these types of clothing, causing a distribution discrepancy between the training and inference stages, as illustrated in Figure 2. Such discrepancies lead to degraded results during the inference stage, where users typically wear tight-fitting garments to facilitate accurate human body tracking.
@@
Our recurrent garment synthesis framework is capable of processing input sequences of arbitrary length while maintaining high efficiency, achieving approximately 10 frames per second on a standard PC. This ensures a smooth and responsive virtual try-on experience for users, as demonstrated in Figure 1. Both qualitative and quantitative evaluations indicate that our method surpasses existing approaches in terms of visual quality and temporal consistency. We also conduct ablation studies to validate the necessity of the garment-invariant representation and the recurrent synthesis framework.
@@ 2.2. Image-based Virtual Try-On
Image-based virtual try-on methods eliminate the need for 3D modeling of garments, relying instead on a single 2D in-shop image to generate try-on results (Song et al., 2023). This makes them significantly more accessible and scalable compared to 3D model-based approaches. Several methods (Jetchev and Bergmann, 2017; Choi et al., 2021; Lee et al., 2022; Kim et al., 2024; Wang et al., 2024; Xu et al., 2024b) focus on producing static try-on images, which lack the ability to convey dynamic garment behavior. To address this, other approaches (Dong et al., 2019; Chen et al., 2021; Jiang et al., 2022; Fang et al., 2024; Xu et al., 2024a; Karras et al., 2024) incorporate inter-frame consistency to generate temporally coherent video try-on results. However, these methods typically suffer from high memory usage, restricting the length of video they can process. It is also important to note that most existing image-based virtual try-on systems are designed primarily for e-commerce retailers to showcase garments on idealized fashion models for marketing purposes. Their networks are typically trained on datasets biased toward tall and slim body types, resulting in suboptimal performance when applied to average users.
In contrast, our method is designed for customer-centric virtual try-ons and supports real-time virtual try-on with temporal consistency for arbitrary lengths of input frames.
@@ 2.4. Domain Adaption
Domain adaptation aims to ensure the performance of machine learning models trained on a source domain when applied to a target domain with a different data distribution (Farahani et al., 2021). The predominant paradigm of domain adaptation is mapping the source and target domains into a shared space (Tzeng et al., 2019, 2017, 2015; Sankaranarayanan et al., 2018; Hoffman et al., 2016). (Hoffman et al., 2016) present a domain adaptation framework with fully convolutional networks for semantic segmentation. They learn the mapping to the shared space in an unsupervised manner by assuming source and target domains share the same label statistics. (Isola et al., 2017) directly learns the mapping from the source domain to the target domain in a supervised way. However, the correlations between the source and target domain are usually unknown. (Royer et al., 2020) and (Zhu et al., 2017) use cycle consistency to learn the mappings between unpaired source and target images. However, unpaired image-to-image translation struggles with high-resolution images, limiting their applicability in our scenario.
@@ 3.1. Overview
Figure 3 provides an overview of our method. To mitigate the distribution gap between the training and inference stages, we aim to obtain robust human body semantic maps under loose-fitting garments. To achieve this, we first train a per-person BodyMap network using images of a person wearing arbitrary tight-fitting garments. This network can then be used to robustly estimate body semantic maps for the same person in the target loose-fitting garment by transforming the extracted garment-invariant representations into body semantic maps.
Subsequently, these estimated maps are used to train a per-garment Recurrent Garment Synthesis (ReGarSyn) network that learns the mapping from the input sequence to the target garment image sequence. Once trained, the ReGarSyn network enables real-time, temporally consistent virtual try-ons for arbitrary persons.
@@ 3.2. Design of Garment-Invariant Representation
(Wu et al., 2025) employ DensePose (Güler et al., 2018) for estimating the human body semantic map directly from the raw input image. However, DensePose varies significantly under loose-fitting garments, as shown in Figure 4. To ensure reliable human body semantic map estimation under loose-fitting garments, we need a garment-invariant representation that captures human body information and is invariant to the garment to serve as the intermediate representation. A natural candidate for such a representation is the set of human joints. However, not all joints are truly garment-invariant. The presence and type of garments can influence their detected positions and associated confidence scores, as shown in Figure 12. The human joint estimation results are generated by Sapiens (Khirodkar et al., 2024), which is the state-of-the-art human pose estimation method.
Although human joint estimation is not entirely garment-invariant, it can still serve as a robust representation by excluding joints sensitive to garment variations, such as the hips and knees. However, excluding these joints limits the expressiveness of the resulting pose representation. To address this limitation, we augment the remaining joint data with virtual measurement garment (Wu et al., 2025), which is a rendered image of the SMPL mesh that encapsulates the full 3D human pose. In contrast to (Wu et al., 2025), which uses an upper-body template mesh, we utilize a full-body template mesh to facilitate the representation of full-body pose.
As illustrated in Figure 5, given an input person image $\mathbf{I}\in\mathbb{R}^{3\times H\times W}$, we first estimate the 3D human pose using BEV (Sun et al., 2022), which outputs a SMPL mesh (Loper et al., 2015). To ensure garment invariance, we fix the body shape parameters that are sensitive to garment variations. Subsequently, we remove non-essential body parts (head, hands, and feet) and apply a grid-pattern texture to the remaining mesh. This textured mesh is then rendered into an RGB image $\mathbf{I}_{vm}\in\mathbb{R}^{3\times H\times W}$.
In parallel, we extract 133 joint heatmaps using Sapiens (Khirodkar et al., 2024). To further reduce sensitivity to garment variation, we discard the heatmaps corresponding to the hip and knee joints. The remaining heatmaps are grouped into three categories and compactly represented as an RGB image $\mathbf{I}_{shm}\in\mathbb{R}^{3\times H\times W}$.
@@ 3.3. Reliable Human Body Semantic Map Estimation
Figure 6. Visual illustration of human body semantic maps: (a) The input image; (b) The DensePose map; (c) The human body semantic map of (Wu et al., 2025); (d) The human body semantic map of ours.
@@
As illustrated in Figure 3, we use the recorded video to generate paired garment-invariant representations and human body semantic maps to train a BodyMap network, whose architecture follows pix2pixHD (Wang et al., 2018). The BodyMap network is then used to estimate the human body semantic map of the same person wearing a loose-fitting garment.
@@ 3.4. Dataset Generation for Loose-Fitting Garments
For each garment item, we collect a garment-specific dataset, hereafter referred to as a per-garment dataset, to enable the training of garment-specific networks. Similar to the procedure in (Wu et al., 2025), we record a video of the human model wearing the loose-fitting garment performing predefined movements. In addition, we record a video of the same human model wearing an arbitrary tight-fitting garment and performing the same predefined movements, which is used to train the BodyMap network.
As illustrated in Figure 7, given a frame $\mathbf{I}$ of the loose-fitting garment videos, we obtain the garment image $\mathbf{I}_{g}$ and garment mask $\mathbf{I}_{m}$ using SAM2 (Ravi et al., 2024). Then, we obtain the virtual measurement garment image $\mathbf{I}_{vm}$, which is also used to construct the garment-invariant representation $\mathbf{I}_{gi}$. By inputting the garment-invariant representation into the trained BodyMap network, we obtain the human body semantic map $\mathbf{I}_{sdp}$ robustly under loose-fitting garments.
Figure 7. Per-garment dataset generation from recorded video frames of the human model. Unlike (Wu et al., 2025) that directly obtain the human body semantic map using DensePose (Güler et al., 2018), we extract garment-invariant representation and then transform it into the semantic map by our trained BodyMap network (red dashed box).
@@ 3.5. Training Pipeline of Virtual Try-On
We follow (Wu et al., 2025) to utilize hybrid person representation as the intermediate representation for garment synthesis. The hybrid person representation $\mathbf{I}_{hybrid}\in\mathbb{R}^{6\times H\times W}$ is a combination of virtual measurement garment $\mathbf{I}_{vm}$ and human body semantic map $\mathbf{I}_{sdp}$, and can be formally represented as:
@@ 3.6. Inference Pipeline of Virtual Try-On
Our recurrent framework facilitates temporally coherent virtual try-on outputs over arbitrarily long input sequences while maintaining real-time performance and a constant memory footprint.
@@ 4.1. Experiment Setting
Implementation details. We set the resolution of the BodyMap network to $512\times 384$ and the ReGarSyn networks to $576\times 432$. For each garment, we trained a per-garment model using the corresponding per-garment dataset. We trained each model on an NVIDIA RTX 4090 GPU for $40$ epochs, with a learning rate of $2\times 10^{-4}$. During training, we randomly sampled clips ranging in length from 8 to 60 frames. We use the Adam optimizer with $\beta_{1}=0.5$ and $\beta_{2}=0.999$. The training duration for a garment-specific model was approximately $12$ hours.
@@ 4.5. Ablation Study on Garment-Invariant Representation
  • VM: Using virtual measurement garment (Wu et al., 2024) as the intermediate representation (Figure 9(c)). This demonstrates the importance of joint heatmaps.

### v3

@@ Abstract.
Per-garment virtual try-on methods collect garment-specific datasets and train networks tailored to each garment to achieve superior results. However, these approaches often struggle with loose-fitting garments due to two key limitations: (1) They rely on human body semantic maps to align garments with the body, but these maps become unreliable when body contours are obscured by loose-fitting garments, resulting in degraded outcomes; (2) They train garment synthesis networks on a per-frame basis without utilizing temporal information, leading to noticeable jittering artifacts. To address the first limitation, we propose a two-stage approach for robust semantic map estimation. First, we extract a garment-invariant representation from the raw input image. This representation is then passed through an auxiliary network to estimate the semantic map. This enhances the robustness of semantic map estimation under loose-fitting garments during garment-specific dataset generation. To address the second limitation, we introduce a recurrent garment synthesis framework that incorporates temporal dependencies to improve frame-to-frame coherence while maintaining real-time performance. We conducted qualitative and quantitative evaluations to demonstrate that our method outperforms existing approaches in both image quality and temporal coherence. Ablation studies further validate the effectiveness of the garment-invariant representation and the recurrent synthesis framework.
Virtual Try-On, Image Synthesis
copyright: noneconference: ; ; $c$s: Computing methodologies Image processingccs: Human-centered computing Mixed / augmented realityccs: Computing methodologies Image-based rendering
@@
Figure 1. (a) Our virtual try-on system features an efficient inference pipeline that enables real-time performance. (b) Our method produces highly realistic $360^{\circ}$ virtual try-on results for loose-fitting garments.
@@ 1. Introduction
Virtual try-on technology, which enables customers to try on garments without physical access, has recently garnered significant attention from researchers in today’s e-commerce landscape. Existing image-based virtual try-on methods (Jetchev and Bergmann, 2017; Dong et al., 2019; Choi et al., 2021; Lee et al., 2022; Fang et al., 2024; Karras et al., 2024) require only a 2D in-shop image of the target garment to generate virtual try-on results, making them highly accessible. However, these approaches are predominantly tailored for retailers aiming to produce promotional images featuring fashion models, rather than for end-users seeking personalized virtual try-on experiences. This focus is clearly reflected in the biased datasets used for training, which are heavily skewed toward tall and slim fashion models. As a result, these methods fail to generalize well to the diverse body types of average consumers.
Unlike general image-based methods, per-garment virtual try-on methods (Chong et al., 2021; Wu et al., 2024, 2025) focus on customer-oriented virtual try-on by collecting garment-specific datasets and training garment-specific neural networks. Early per-garment methods (Chong et al., 2021; Wu et al., 2024) require an expensive robotic mannequin for dataset collection, making their results non-reproducible. A follow-up work (Wu et al., 2025) lowers the barrier to dataset collection significantly by using real human bodies for dataset collection. This approach relies on DensePose (Güler et al., 2018) for estimating human body semantic maps, which are crucial for aligning the human body and the synthesized garment in 2D image space. However, it struggles with loose-fitting garments, as DensePose results degrade significantly for these types of clothing, causing a distribution discrepancy between the training and inference stages, as illustrated in Figure 2. Such discrepancies lead to degraded results during the inference stage, where users typically wear tight-fitting garments to facilitate accurate human body tracking.
@@
Our recurrent garment synthesis framework is capable of processing input sequences of arbitrary length while maintaining high efficiency, achieving approximately 10 frames per second on a standard PC. This ensures a smooth and responsive virtual try-on experience for users, as demonstrated in Figure 1. We performed both qualitative and quantitative comparisons against general image-based methods and the existing per-garment method to demonstrate the superiority of our approach in terms of visual quality and temporal consistency. In addition, we validated the effectiveness of the garment-invariant representation by replacing it with alternative representations and assessed the necessity of the recurrent synthesis framework through ablation studies.
@@ 2.2. Image-based Virtual Try-On
Image-based virtual try-on methods eliminate the need for 3D modeling of garments, relying instead on a single 2D in-shop image to generate try-on results (Song et al., 2023). This makes them significantly more accessible and scalable compared to 3D model-based approaches. Several methods (Jetchev and Bergmann, 2017; Choi et al., 2021; Lee et al., 2022; Kim et al., 2024; Wang et al., 2024; Xu et al., 2024b) focus on producing static try-on images, which lack the ability to convey dynamic garment behavior. To address this, other approaches (Dong et al., 2019; Chen et al., 2021; Jiang et al., 2022; Fang et al., 2024; Xu et al., 2024a; Karras et al., 2024) incorporate inter-frame consistency to generate temporally coherent video try-on results. However, these methods typically suffer from high memory usage, restricting the length of video they can process. In addition, most image-based methods rely on datasets scraped from online images featuring tall and slim fashion models, which leads to suboptimal performance when applied to average users.
In contrast, our method enables real-time virtual try-on with temporal consistency across arbitrary video lengths and demonstrates strong generalization to unseen body shapes.
@@ 2.4. Domain Adaption
Domain adaptation aims to ensure the performance of machine learning models trained on a source domain when applied to a target domain with a different data distribution (Farahani et al., 2021). The predominant paradigm of domain adaptation is mapping the source and target domains into a shared space (Tzeng et al., 2019, 2017, 2015; Sankaranarayanan et al., 2018; Hoffman et al., 2016). Hoffman et al. (Hoffman et al., 2016) present a domain adaptation framework with fully convolutional networks for semantic segmentation. They learn the mapping to the shared space in an unsupervised manner by assuming source and target domains share the same label statistics. Isola et al. (Isola et al., 2017) directly learns the mapping from the source domain to the target domain in a supervised way. However, the correlations between the source and target domains are usually unknown. The works in (Royer et al., 2020) and (Zhu et al., 2017) use cycle consistency to learn the mappings between unpaired source and target images. However, unpaired image-to-image translation struggles with high-resolution images, limiting their applicability in our scenario.
@@ 3.1. Overview
Figure 3 provides an overview of our method. To mitigate the distribution gap between the training and inference stages, we aim to obtain robust human body semantic maps under loose-fitting garments. To achieve this, we propose a garment-invariant representation that is robust to garment variations. Using images of a person wearing a tight-fitting garment, we train a per-person BodyMap network to translate this representation into the body semantic map.
Once trained, the BodyMap network can be used to robustly estimate body semantic maps for the same person in the target loose-fitting garment by transforming the extracted garment-invariant representations into body semantic maps. These estimated maps are then used to train a per-garment Recurrent Garment Synthesis (ReGarSyn) network that learns the mapping from the input sequence to the target garment image sequence.
Unlike prior works (Chong et al., 2021; Wu et al., 2024, 2025), which synthesize garment images frame by frame, our proposed ReGarSyn network synthesizes garment images in a recurrent manner to enhance temporal consistency. During inference, our recurrent framework facilitates temporally coherent virtual try-on for arbitrary persons over long input sequences while maintaining real-time performance and a constant memory footprint.
While it is technically feasible to use the garment-invariant representation as input to the ReGarSyn network, which eliminates the need for the BodyMap network, we opted against this approach due to concerns regarding inference speed. Generating the garment-invariant representation is computationally intensive, whereas the semantic map can be extracted in real-time. Therefore, it is necessary to train an auxiliary BodyMap network that maps the garment-invariant representation to the semantic map, thereby allowing us to bypass the computational overhead of generating the garment-invariant representation during inference.
@@ 3.2. Design of Garment-Invariant Representation
The prior work (Wu et al., 2025) employs DensePose (Güler et al., 2018) for estimating the human body semantic map directly from the raw input image. However, DensePose varies significantly under loose-fitting garments, as shown in Figure 4. To ensure reliable human body semantic map estimation under loose-fitting garments, we need a garment-invariant representation that captures human body information and is invariant to the garment to serve as the intermediate representation. A natural candidate for such a representation is the set of human joints. However, not all joints are truly garment-invariant. The presence and type of garments can influence their detected positions and associated confidence scores, as shown in Figure 5. The human joint estimation results are generated by Sapiens (Khirodkar et al., 2024), which is the state-of-the-art human pose estimation method.
((a)) Garment’s influence on the human joint locations.
((b)) Garment’s influence on human joint confidence scores.
Figure 5. Human joint estimation is not fully garment-invariant, especially for the knee and hip joints, whose positions and confidence scores are sensitive to garment variations.
Although human joint estimation is not entirely garment-invariant, it can still serve as a robust representation by excluding joints sensitive to garment variations, such as the hips and knees. However, excluding these joints limits the expressiveness of the resulting pose representation. To address this limitation, we augment the remaining joint data with a virtual measurement garment (Wu et al., 2025), which is a rendered image of the SMPL mesh that encapsulates the full 3D human pose. In contrast to (Wu et al., 2025), which uses an upper-body template mesh, we utilize a full-body template mesh to facilitate the representation of full-body pose.
As illustrated in Figure 6, given an input person image $\mathbf{I}\in\mathbb{R}^{3\times H\times W}$, where $H$ and $W$ denote the height and width respectively, we first estimate the 3D human pose using BEV (Sun et al., 2022), which outputs a SMPL mesh (Loper et al., 2015). To ensure garment invariance, we fix the body shape parameters $\mathbf{\beta}\in\mathbb{R}^{10}$, which are sensitive to garment variations, to a zero vector. Subsequently, we remove non-essential body parts (head, hands, and feet) and apply a grid-pattern texture proposed by (Halimi et al., 2022) to the remaining mesh, which enhances the expressiveness of human body orientation. This textured mesh is then rendered into an RGB image $\mathbf{I}_{vm}\in\mathbb{R}^{3\times H\times W}$.
In parallel, we extract 133 joint heatmaps using Sapiens (Khirodkar et al., 2024). To further reduce sensitivity to garment variation, we discard the heatmaps corresponding to the hip and knee joints. The remaining heatmaps are grouped into three categories: left-side limb joints, right-side limb joints, and head joints. These are compactly represented as a three-channel RGB image, denoted as $\mathbf{I}_{shm}\in\mathbb{R}^{3\times H\times W}$. As illustrated in Figure 6, the R channel represents right-side limb joints, the G channel encodes head joints, and the B channel corresponds to left-side limb joints.
@@ 3.3. Reliable Human Body Semantic Map Estimation
Figure 7. Visual illustration of human body semantic maps: (a) The input image; (b) The DensePose map; (c) The human body semantic map of (Wu et al., 2025); (d) Our proposed human body semantic map.
Figure 8. The boundary between the lower and upper body of the semantic map generated by DensePose (Güler et al., 2018) is sensitive to garment variation, even for tight-fitting garments.
@@
As illustrated in Figure 3, we use the recorded video to generate paired garment-invariant representations and human body semantic maps to train the BodyMap network, whose architecture follows pix2pixHD (Wang et al., 2018). The BodyMap network is then used to estimate the human body semantic map of the same person wearing a loose-fitting garment.
@@ 3.4. Dataset Generation for Loose-Fitting Garments
For each garment item, we collect a garment-specific dataset, hereafter referred to as a per-garment dataset, to enable the training of garment-specific networks. Similar to the procedure in (Wu et al., 2025), we record a video of the human model wearing the loose-fitting garment performing predefined movements. In addition, we record a separate video of the same human model wearing a tight-fitting garment and performing the same predefined movements, which is used to train the BodyMap network.
As illustrated in Figure 9, given a frame $\mathbf{I}$ of the loose-fitting garment videos, we extract the garment image $\mathbf{I}_{g}$ and garment mask $\mathbf{I}_{m}$ using SAM2 (Ravi et al., 2024). In addition, we generate the virtual measurement garment image $\mathbf{I}_{vm}$, which is also used to construct the garment-invariant representation $\mathbf{I}_{gi}$. By inputting the garment-invariant representation into the trained BodyMap network, we obtain the human body semantic map $\mathbf{I}_{sdp}$ robustly under loose-fitting garments.
Figure 9. Per-garment dataset generation from recorded video frames of the human model. Unlike (Wu et al., 2025) that directly obtains the human body semantic map using DensePose (Güler et al., 2018), we extract a garment-invariant representation and then transform it into the semantic map by our trained BodyMap network (red dashed box).
@@ 3.5. Training Pipeline of Virtual Try-On
We follow (Wu et al., 2025) to utilize a hybrid person representation as the intermediate representation for garment synthesis. The hybrid person representation $\mathbf{I}_{hybrid}\in\mathbb{R}^{6\times H\times W}$ is a combination of virtual measurement garment $\mathbf{I}_{vm}$ and human body semantic map $\mathbf{I}_{sdp}$, and can be formally represented as:
@@
Figure 10. Detailed illustration of the ReGarSyn network training process.
@@ 3.6. Inference Pipeline of Virtual Try-On
Figure 11. Detailed illustration of the inference pipeline for the ReGarSyn network.
@@ 4.1. Experiment Setting
Implementation details. We set the resolution of the BodyMap network to $512\times 384$ and the ReGarSyn networks to $576\times 432$. For each garment, we trained a per-garment model using the corresponding per-garment dataset. We trained each model on an NVIDIA RTX 4090 GPU for $40$ epochs, with a learning rate of $2\times 10^{-4}$. During training, we randomly sampled clips ranging in length from 8 to 60 frames. We use the Adam optimizer with $\beta_{1}=0.5$ and $\beta_{2}=0.999$. The training duration for a garment-specific network was approximately $12$ hours.
@@ 4.2. Qualitative Evaluation
Figure 13. Virtual try-on results for unseen body shapes. Our ReGarSyn networks trained on per-garment datasets collected by a single human model demonstrate strong generalization to unseen body shapes.
@@ 4.5. Ablation Study on Garment-Invariant Representation
  • VM: Using a virtual measurement garment (Wu et al., 2024) as the intermediate representation (Figure 14(c)). This demonstrates the importance of joint heatmaps.
@@
Figure 15. Ablation study on garment-invariant representation. DP: The semantic map estimated by DensePose (Güler et al., 2018) is not robust against loose-fitting garments. HM: Using joint heatmaps that include garment-sensitive joints leads to the inflation of the semantic map. SHM: Using joint heatmaps without garment-sensitive joints fails to provide adequate information for semantic map estimation. VM: Using solely virtual measurement garments fails to provide pixel-level accurate guidance for semantic maps.